Western Industrial Products is considering a project with a four-year life and an initial cost of $268,000. The discount rate for the project is 14.5 percent. The firm expects to sell 2,200 units on the last day of each year. The cash flow per unit is $50. The firm will have the option to abandon this project after one year at which time the project's assets could be sold for an estimated $225,000. The firm should abandon the project immediately following the sale on the last day of the first year if the expected level of annual sales, starting with year 2, falls to _____ units or less. Ignore taxes. A. 1,955 units B. 1,667 units C. 1,922 units D. 2,034 units E. 2,108 units
Answer #1

Present value of cash flow must be equals to sale value of assets at the end of year 1 from abandonment of project in order to breakeven.

Present value of cash flow from year 1-3 =PVA14.5%,3* Cash flow per unit *Number of units

225000 = 2.30229*50*number of units

225000 = 115.1145 *number of units

Number of units = 225000/115.1145

                           = 1954.57   (rounded to 1955)

The firm should abandon the project immediately following the sale on the last day of the first year if the expected level of annual sales, starting with year 2, falls to 1955 or less /.

correct option is "A"

**Find present value annuity factor using financial calculator where i =14.5% ,n= 3   (since out of 4 years ,1 year is expired) PMT = 1